My problem is this:
Solving by elimination:
9x+4y = -16
3x-4y = -8
I have worked it out several times and don't get it.
9x+4y = -16
3x-4y = -8
Rewrittten:
12x = -24
x=-2
Rewritten:
9(-2)+4y=-16
4y= -16 9(-2)
y=-34
I then divide -34/4 and get 8 1/2. I still keep getting it wrong on my online homework. Please advise me as to what I am doing wrong.

Answer by lwsshak3(11628)

My problem is this:
Solving by elimination:
9x+4y = -16
3x-4y = -8
add
12x=-24
x=-2
4y=-16-9x=-16-18=-34
y=-34/4=-8.5
...
Your answer is correct except for the sign. Answer should be negative.